OpenStudy (anonymous):

How do i graph y=3/x?

OpenStudy (anonymous):

Well an easy way would be to use random numbers, and see the output. So like y = 3/1 = 3 so your first point would be (1,3) ^.^, easy way to do it :3, it'll be an asymptote like feature :D

OpenStudy (anonymous):

Because 3/0 = undefined

OpenStudy (the_fizicx99):

A simple way to do it is by making a table of, -1,-2,0,1,2 Sub in those values for x and plot them

OpenStudy (anonymous):